# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ Appendix F Additional Details on Epsilon ($\epsilon$)
Figure 4: Ablation on epsilon values across multiple datasets.
Appendix G Detailed benefits of our multi-template distillation approach.
The proposed multi-template distillation strategy resembles to the unsupervised approach employed to leverage pseudo-labels in semi-supervised learning, e.g., [2]. In this learning scenario, unlabeled samples are augmented, whose predictions are averaged to obtain a pseudo-label as $\mathbf{q}_{u}=\frac{1}{M}\sum_{m}f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$, where $\bar{\mathbf{x}}_{u,m}$ is the $m$-th augmented version of a given unlabeled image $\mathbf{x}_{u}$. Once the average pseudo-label is obtained, we can minimize the pseudo cross-entropy as $-\sum_{k}\mathbf{q}_{u}\log f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$ [^3] For simplicity, we consider only one unlabeled image.. In this setting, the predictions from multiple augmented images are guided by the average pseudo-label. Once the loss is computed over all the predictions from the augmented images, the gradient is back-propagated and the network parameters updated. Limitations of this method: First, this approach forces all augmentations to align with a single static target $\mathbf{q}$, potentially causing the model to converge to a suboptimal solution because it cannot make incremental adjustments based on each individual augmentation. Second, it treats all augmentations as being equally aligned with a single target $\mathbf{q}$, which could limit the model’s ability to handle diverse augmentations, especially if some augmentations are very different from others. And last, this strategy applies a global update that may force the model too quickly into a solution where all augmentations have the same prediction target $\mathbf{q}$, potentially making it harder for the model to learn robust features. To overcome these limitations in the TTA scenario, where no single supervision is available, the proposed method takes an opposite direction to these well-established practices. More concretely, our approach minimizes the following pseudo cross-entropy loss at each step $m$: $-\mathbf{q}_{m}\log\mathbf{p}$, where $\mathbf{q}_{m}$ is the pseudo-label at step $m$ (derived from the $m$-th text prompt, which is equivalent to one augmentation), and $\mathbf{p}$ is the mean prediction considering all the $M$ text prompts (similar to averaging over $M$ augmentations). At step $m$, the gradient is backpropagated, the network parameters updated. Then, we can recompute the pseudo-labels at step $m+1$, as well as the new mean prediction $\mathbf{p}$ (which will be now different due to the parameters update). Advantages: First, by updating the model after each step and recomputing $\mathbf{p}$ as the mean of the model’s outputs, our approach encourages a more fine-grained, iterative refinement of predictions, potentially helping the model to converge more smoothly. Indeed, the model can adjust its predictions for each template (i.e., augmentation) before being evaluated again in the next step. Furthermore, this step-by-step refinement process allows the model to better adapt to the diversity in the augmentations and learn a more generalizable representation.
Appendix H Discussion on the Two Hyperparameters of WATT and Motivation for CLIP-OT
WATT [31] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process). While these hyperparameters play an essential role in improving WATT’s adaptation capabilities, they introduce significant computational and scalability challenges, particularly as the number of classes grows. The iterative nature of $L$ requires multiple forward and backward passes for every template, leading to substantial runtime overhead, especially for datasets with many classes like Tiny-ImageNet. Additionally, higher $L$ values, while effective for complex corruptions, risk overfitting to noisy pseudo-labels, while lower $L$ values may lead to under-adaptation. Similarly, $M$ stabilizes adaptation through repeated weight averaging, but its effectiveness diminishes beyond a certain point, with larger values providing negligible performance improvements while exponentially increasing runtime. This dependence on repeated updates for each class and template becomes especially prohibitive for datasets with a large number of classes, as demonstrated in Figure 3, where runtime increases from CIFAR-10 (10 classes) to Tiny-ImageNet (200 classes).
These limitations motivate our proposed approach, CLIP-OT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$. CLIP-OT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging. This not only drastically reduces computational overhead but also ensures scalable, robust adaptation across diverse corruption levels and datasets, offering a practical alternative to WATT’s hyperparameter-dependent framework.
Appendix I Further results on other backbones
In this section, we present additional experimental results leveraging different backbone architectures to evaluate the robustness and generalizability of our proposed method. Specifically, we analyze the performance of CLIP models with ViT-B/16 and ViT-L/14 backbones across various datasets, including CIFAR, Tiny-ImageNet and their corruptions. These experiments validate our approach across diverse settings and highlight its adaptability to challenging domain shifts.
Table 12 details the performance of CLIP-OT with the ViT-B/16 backbone on CIFAR-10, CIFAR-10.1, and CIFAR-10-C datasets. Our proposed method demonstrates consistent improvements across both clean and corrupted data, outperforming baseline methods such as CLIP, TENT, and TPT. Similarly for ViT-L/14 backbone on these datasets, we obtain similar gains as illustrated in Table 13. Table 14 extends the evaluation to CIFAR-100 and CIFAR-100-C, demonstrating the effectiveness of our method on more complex datasets with larger label spaces. Using CLIP with the ViT-B/16 backbone, our approach achieves the highest average accuracy, with substantial improvements under corruptions such as Glass Blur ($+5.30\%$) and Contrast ($+6.23\%$). We also evaluate these datasets with backbone ViT/L-14 in Table 15.
For Tiny-ImageNet and its corrupted version, Tiny-ImageNet-C, Table 16 and Table 17 provides a detailed breakdown. On datasets with higher complexity and larger label spaces, such as Tiny-ImageNet and Tiny-ImageNet-C (Table 16), CLIP-OT exhibits substantial gains, consistent with the main paper’s observations. Specifically, our method improves over WATT-S by over 6% on average across Tiny-ImageNet-C corruptions, showcasing its scalability to more challenging scenarios.
I.1 Tiny-Imagenet Experiments
We report the mean and standard deviation over Tiny-Imagenet experiments (Table 2, main paper) in Table 18.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05 92.31 $\pm$0.10 91.97 $\pm$0.03 94.69 $\pm$0.18
CIFAR-10.1 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33 87.90 $\pm$0.11 88.10 $\pm$0.08 89.25 $\pm$0.08
Gaussian Noise 37.75 31.04 $\pm$0.38 35.35 $\pm$0.15 60.89 $\pm$0.26 63.10 $\pm$0.12 65.57 $\pm$0.22 68.50 $\pm$0.19
Shot Noise 41.10 40.54 $\pm$0.41 41.03 $\pm$0.19 65.19 $\pm$0.21 66.31 $\pm$0.10 68.67 $\pm$0.37 71.24 $\pm$0.22
Impulse Noise 51.71 58.03 $\pm$0.16 54.86 $\pm$0.07 67.55 $\pm$0.09 69.62 $\pm$0.12 70.39 $\pm$0.11 76.97 $\pm$0.17
Defocus Blur 70.07 77.57 $\pm$0.03 70.29 $\pm$0.02 78.92 $\pm$0.12 79.64 $\pm$0.08 79.90 $\pm$0.07 83.33 $\pm$0.09
Glass Blur 42.24 47.16 $\pm$0.05 37.86 $\pm$0.17 57.18 $\pm$0.20 58.98 $\pm$0.12 61.62 $\pm$0.21 68.33 $\pm$0.20
Motion Blur 65.81 76.16 $\pm$0.05 67.43 $\pm$0.11 76.59 $\pm$0.06 78.32 $\pm$0.16 79.02 $\pm$0.07 83.13 $\pm$0.10
Zoom Blur 72.50 79.64 $\pm$0.12 72.91 $\pm$0.02 79.62 $\pm$0.11 80.67 $\pm$0.07 81.10 $\pm$0.08 85.35 $\pm$0.11
Snow 73.23 81.68 $\pm$0.03 72.98 $\pm$0.32 81.13 $\pm$0.29 81.99 $\pm$0.10 82.54 $\pm$0.18 86.65 $\pm$0.12
Frost 76.52 83.22 $\pm$0.05 75.87 $\pm$0.16 81.24 $\pm$0.08 83.41 $\pm$0.16 83.46 $\pm$0.15 86.37 $\pm$0.09
Fog 68.35 80.78 $\pm$0.15 69.13 $\pm$0.27 78.47 $\pm$0.19 81.36 $\pm$0.12 81.88 $\pm$0.12 85.60 $\pm$0.24
Brightness 83.36 89.85 $\pm$0.11 83.67 $\pm$0.14 88.66 $\pm$0.15 89.06 $\pm$0.05 89.10 $\pm$0.14 92.64 $\pm$0.11
Contrast 61.90 79.24 $\pm$0.19 62.16 $\pm$0.06 75.15 $\pm$0.07 81.57 $\pm$0.23 83.79 $\pm$0.12 86.92 $\pm$0.18
Elastic Transform 53.16 62.54 $\pm$0.08 51.26 $\pm$0.23 69.49 $\pm$0.08 69.14 $\pm$0.09 70.93 $\pm$0.20 75.18 $\pm$0.14
Pixelate 48.48 67.08 $\pm$0.24 44.65 $\pm$0.21 71.80 $\pm$0.16 73.38 $\pm$0.29 75.67 $\pm$0.32 78.73 $\pm$0.30
JPEG Compression 56.05 65.42 $\pm$0.05 56.73 $\pm$0.07 66.42 $\pm$0.25 69.02 $\pm$0.10 69.65 $\pm$0.23 72.65 $\pm$0.18
Mean 60.15 68.00 59.75 73.22 75.04 76.22 80.11

### v3

@@ B.3 Additional Details on Epsilon ($\epsilon$)
B.4 Sensitivity to Batch Size
Our method, SAT, is fundamentally a batch-aware approach. We evaluate SAT’s performance on CIFAR-10C while varying the batch size from 16 to 128, comparing it against other methods. As shown in Table 9, SAT’s performance is robust and scales well, achieving strong results even with a batch size of 32. Notably, under distribution shift on CIFAR-10C, SATexhibits substantially lower performance degradation when reducing the batch size (only –2.5%, from 77.35 to 74.40) compared to much larger drops observed for WATT (–7.5%) and BATCLIP (–6%), highlighting SAT’s superior robustness to batch size constraints in corrupted settings. Performance for all methods generally improves with a larger batch, confirming that a sufficiently large batch is important for the stability of TTA.
Table 9: Accuracy (%) on CIFAR-10 and CIFAR-10C vs. Batch Size. Results are on ViT-B/32.
Dataset Method 16 32 64 128
CIFAR-10 WATT 89.14 89.51 90.16 91.05
BATCLIP 84.31 85.22 88.34 90.35
SAT (Ours) 87.06 92.34 93.05 93.22
CIFAR-10C WATT 65.66 68.34 71.21 73.00
BATCLIP 63.08 61.89 65.51 69.61
SAT (Ours) 74.40 77.01 77.35 77.06
B.5 On the impact of multiple templates.
We now assess how the performance evolves as the number of templates increases (Table 10). Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$. For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset. Results from this experiment show that SAT performance consistently increases with the number of templates. This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
Table 10: Accuracy (%) of SAT with varying numbers of templates (see Table 5) on natural domain shift benchmarks.
#\# Templates 1 2 3 4 5 6 7 8
CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
C Detailed Experimental Results
To verify that our adaptation method does not simply overfit to corruptions at the cost of performance on the original data distribution, we evaluate SAT on clean, non-corrupted benchmarks. Table 11 presents the performance of SAT and competing baselines on the standard CIFAR-10, CIFAR-10.1, CIFAR-100, and Tiny-ImageNet test sets.
C.1 Performance on Clean Datasets
Table 11 details the performance of all ViT-B/32 methods on the original, uncorrupted test sets. These results correspond to the clean dataset performance reported in the ablations (Figure 3) and form the basis for the corruption-free evaluation. The results show that SAT does not just preserve performance but consistently improves accuracy over the CLIP baseline and other TTA methods, even without a significant domain shift. On CIFAR-10, SAT achieves $93.15\%$ accuracy, surpassing WATT at $91.05\%$. This trend is even more pronounced in more complex datasets. On CIFAR-100, SAT achieves $71.68\%$, a full 10-point gain over the $61.68\%$ from CLIP and is notably higher than all other baselines. Similarly, on Tiny-ImageNet, SAT $63.69\%$ provides a clear improvement over both the baseline ($58.29\%$) and WATT ($61.35\%$).
Table 11: Performance on Clean Datasets (ViT-B/32). Best method in bold.
Dataset CLIP TENT TPT CLIPArTT WATT BATCLIP SAT
CIFAR-10 88.74 91.69 88.06 90.04 91.05 88.89 93.15
CIFAR-10.1 83.25 87.60 81.80 86.35 86.98 83.70 88.37
CIFAR-100 61.68 69.74 63.78 69.79 70.74 63.94 71.68
Tiny-IN 58.29 57.72 58.90 59.85 61.35 58.93 63.69
C.2 Per-Corruption Results
To supplement the mean accuracy results presented in the main paper (Table 1), this section provides a granular, per-corruption performance breakdown. This analysis demonstrates that SAT’s strong average performance stems from a consistent robustness across a wide variety of distribution shifts.
We extend this detailed analysis to the larger-scale backbones, ViT-B/16 (Table 12) and ViT-L/14 (Table 13).
On ViT-B/16 (Table 12): SAT achieves the highest accuracy on all 15 corruption types for both CIFAR-10C and CIFAR-100C. For example, on CIFAR-10C Pixelate corruption, SAT ($78.73\%$) significantly outperforms the baseline, WATT ($75.67\%$). The gains are similarly strong on ImageNet-C, where SAT (mean $28.35\%$) consistently outperforms all other methods.
On ViT-L/14 (13): The performance gap widens further, demonstrating SAT’s superior scalability. On CIFAR-100C, SAT (mean $69.48\%$) again dominates every single corruption type. The improvement is validated on challenging corruptions like Glass Blur, where SAT ($48.19\%$) more than doubles the performance of the zero-shot baseline ($23.46\%$) and is far ahead of the next-best method, WATT ($33.54\%$). On the difficult Tiny-ImageNet-C, SAT’s mean accuracy of $50.36\%$ is a major leap from the $34.98\%$ baseline and all other TTA methods.
This comprehensive, per-corruption breakdown confirms that SAT provides a fundamentally more robust adaptation mechanism that is effective against diverse types of data shifts, and that these benefits are amplified when applied to larger, more capable vision-language models.
Table 12: Unified corruption robustness results (ViT-B/16, severity = 5). Top-1 accuracy (%) on CIFAR-10C, CIFAR-100C, Tiny-ImageNet-C, and ImageNet-C. Best method in bold (with gray cell), second best underlined.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixelate JPEG Mean $\uparrow$
CIFAR-10C
CLIP 37.75 41.10 51.71 70.07 42.24 65.81 72.50 73.23 76.52 68.35 83.36 61.90 53.16 48.48 56.05 60.15
TENT 31.04 40.54 58.03 77.57 47.16 76.16 79.64 81.68 83.22 80.78 89.85 79.24 62.54 67.08 65.42 68.00
TPT 35.35 41.03 54.86 70.29 37.86 67.43 72.91 72.98 75.87 69.13 83.67 62.16 51.26 44.65 56.73 59.75
WATT 65.57 68.67 70.39 79.90 61.62 79.02 81.10 82.54 83.46 81.88 89.10 83.79 70.93 75.67 69.65 76.22
CLIPArTT 60.89 65.19 67.55 78.92 57.18 76.59 79.62 81.13 81.24 78.47 88.66 75.15 69.49 71.80 66.42 73.22
BATCLIP 61.81 65.49 64.32 80.03 53.82 79.67 81.42 82.63 83.60 80.84 87.20 81.17 69.44 63.17 68.21 73.52
Ours 68.50 71.24 76.97 83.33 68.33 83.13 85.35 86.65 86.37 85.60 92.64 86.92 75.18 78.73 72.65 80.11
CIFAR-100C
CLIP 15.88 17.49 21.43 40.10 13.48 39.82 45.45 42.77 45.39 38.98 52.55 33.32 24.39 21.89 27.21 32.01
TENT 12.28 15.07 13.13 50.35 4.84 49.85 54.76 52.38 51.66 50.74 64.26 48.69 33.56 36.20 30.80 37.90
TPT 15.43 16.88 22.12 41.08 18.43 40.85 46.77 47.24 48.61 39.92 55.83 33.13 27.36 21.26 30.97 33.73
WATT 35.95 37.96 44.62 53.80 33.39 52.72 57.51 56.73 56.48 53.83 66.67 55.06 40.37 47.02 42.13 48.95
CLIPArTT 19.01 20.27 17.66 49.86 18.34 50.00 54.13 52.80 49.56 49.92 63.76 47.86 32.93 39.49 35.56 40.08
BATCLIP 21.79 28.07 30.81 45.46 21.16 45.04 49.82 49.52 47.81 47.32 59.12 43.72 30.89 30.17 32.06 38.85
Ours 38.80 40.89 46.39 55.50 37.89 55.10 58.77 57.84 58.09 55.68 67.76 58.29 44.32 49.46 43.87 51.24
Tiny-ImageNet-C
CLIP 4.80 6.44 4.36 22.23 7.50 33.46 29.86 31.64 33.26 24.26 40.48 1.72 23.85 22.35 27.65 20.92
TENT 13.44 17.46 9.09 32.33 12.47 44.17 39.24 38.91 42.04 31.06 49.26 3.06 34.57 37.73 41.92 29.78
TPT 10.53 12.90 6.57 29.16 10.88 39.91 36.62 38.66 40.95 29.22 47.78 3.58 32.11 29.02 36.51 26.96
WATT 16.78 19.27 10.42 33.33 16.26 44.21 40.26 42.31 44.83 34.08 52.16 5.43 35.11 37.47 42.93 31.66
CLIPArTT 19.25 21.80 13.46 33.49 18.33 45.17 40.53 42.71 44.28 34.78 51.61 7.24 35.16 41.11 44.58 32.90
BATCLIP 11.92 15.36 10.14 31.77 14.80 43.38 39.13 38.92 38.98 31.88 49.11 5.54 32.91 36.72 38.98 29.30
Ours 24.88 27.99 20.69 37.27 22.72 48.75 45.16 45.74 48.09 41.21 55.75 13.06 41.69 45.83 46.55 37.69
ImageNet-C
CLIP 10.70 11.28 10.66 19.60 13.40 18.06 17.54 26.78 27.08 32.86 49.02 14.98 11.94 30.00 29.38 20.89
TENT 7.98 8.68 8.82 21.64 17.08 21.38 19.78 28.26 28.40 33.86 49.54 18.10 13.18 32.56 32.54 22.79
TPT 7.15 6.23 7.03 20.11 11.24 21.36 23.19 26.04 31.17 32.01 50.12 21.05 8.13 39.22 33.15 22.35
WATT 7.68 8.80 9.14 23.70 18.36 24.88 22.54 29.56 29.36 35.52 50.82 21.44 14.60 34.90 34.40 24.38
CLIPArTT 19.10 19.80 19.10 14.10 18.50 19.40 17.00 26.40 29.00 35.20 47.40 2.20 22.20 28.10 34.65 23.47
BATCLIP 17.64 19.00 16.50 24.46 19.04 27.20 24.20 31.90 29.74 37.46 52.20 21.56 18.82 32.12 35.56 27.16
Ours 19.80 21.14 19.20 24.78 21.38 26.82 24.40 32.04 32.36 38.54 49.44 24.42 20.16 35.92 34.84 28.35
Table 13: Unified corruption robustness results (ViT-L/14, severity = 5). Top-1 accuracy (%) on CIFAR-10C, CIFAR-100C, Tiny-ImageNet-C, and ImageNet-C. Best method in bold (with gray cell), second best underlined.
Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixelate JPEG Mean $\uparrow$
CIFAR-10C
CLIP 64.64 67.82 78.21 80.73 50.29 80.75 82.75 83.01 84.90 78.44 91.67 84.20 65.45 75.10 72.58 76.04
TENT 68.87 71.95 80.22 83.10 57.12 82.69 84.91 85.99 87.15 81.30 93.07 87.93 69.96 78.88 75.49 79.18
TPT 64.44 66.81 76.46 79.01 49.64 78.85 82.32 82.69 84.63 77.56 90.94 82.88 64.81 77.89 71.18 75.01
WATT 72.73 74.60 80.95 83.15 62.35 82.61 85.44 85.61 86.88 81.79 92.59 87.38 71.25 77.67 75.84 80.06
CLIPArTT 70.04 71.44 79.42 81.75 58.13 80.76 83.39 84.48 85.21 79.27 91.87 86.19 67.43 77.88 74.46 78.06
BATCLIP 75.04 79.39 84.19 86.93 71.01 86.80 87.56 89.49 89.44 86.88 94.04 92.00 79.42 84.38 80.32 83.79
Ours 80.47 82.64 87.80 90.33 77.72 88.93 91.69 91.56 91.43 90.82 95.86 94.53 82.16 87.13 82.60 86.35
CIFAR-100C
CLIP 30.55 34.58 44.89 48.88 23.46 50.83 56.02 49.03 53.27 48.51 60.53 50.24 35.07 43.86 39.11 44.59
TENT 36.93 38.23 49.09 55.23 27.02 56.03 61.19 55.60 58.21 49.26 67.34 59.91 38.49 48.37 43.44 50.14
TPT 36.10 40.96 49.69 50.43 24.35 51.94 56.96 54.89 58.15 53.37 66.60 53.64 35.72 44.32 44.42 47.58
WATT 44.13 46.63 56.26 57.66 33.54 57.81 62.74 61.04 62.76 54.70 71.60 63.95 41.27 51.22 49.78 54.34
CLIPArTT 41.46 44.27 51.44 56.55 30.47 56.98 62.56 58.81 60.38 54.38 69.63 63.39 39.57 50.45 47.45 52.52
BATCLIP 39.87 43.50 50.11 51.55 30.95 51.03 55.98 53.77 54.53 49.08 65.52 59.07 34.07 47.96 45.61 48.84
Ours 52.26 54.59 63.02 65.20 48.19 65.09 68.92 66.81 67.71 65.54 76.25 72.34 52.72 59.55 54.98 62.21
Tiny-ImageNet-C
CLIP 17.26 21.57 17.32 35.63 13.62 50.11 43.65 48.04 49.36 37.58 55.85 5.78 34.15 45.90 50.13 34.98
TENT 26.60 28.93 22.22 38.66 10.98 52.53 46.02 49.13 50.46 38.63 58.25 8.66 36.81 50.77 51.88 40.28
TPT 23.22 27.39 19.79 39.44 15.72 53.91 48.82 51.50 53.89 40.43 61.07 7.48 40.65 52.28 52.38 41.07
WATT 27.98 31.47 25.38 40.55 17.44 54.46 48.28 52.60 53.77 42.57 62.01 9.21 39.85 51.83 54.19 43.28
CLIPArTT 29.27 33.01 27.93 39.99 17.09 53.63 47.39 51.97 51.49 42.35 60.69 11.25 38.44 52.22 52.46 42.98
BATCLIP 23.98 27.34 23.06 35.40 14.42 48.27 42.72 41.82 46.54 36.29 54.78 6.90 36.84 47.59 49.89 35.72
Ours 38.14 41.02 35.48 46.03 27.93 57.78 52.67 56.85 57.56 50.49 65.28 21.57 47.86 58.23 56.82 50.36
ImageNet-C
CLIP 19.26 21.62 17.88 27.20 20.26 31.78 29.22 43.16 38.08 43.38 61.52 29.50 20.76 44.22 32.94 32.05
TENT 20.48 22.70 15.86 28.00 21.74 33.40 30.44 43.96 38.58 43.92 61.88 32.22 22.22 45.84 35.10 33.09
TPT 19.00 26.00 17.00 28.00 16.00 30.00 30.00 42.00 31.00 41.00 59.00 32.00 15.00 43.00 31.00 30.67
WATT 23.32 26.18 6.06 32.26 27.66 38.54 34.96 47.14 41.52 46.46 63.32 38.66 26.92 50.10 41.38 36.30
CLIPArTT 27.10 29.00 29.20 21.80 21.30 31.00 28.40 39.60 38.50 41.50 64.60 18.40 30.70 43.30 48.00 34.13
BATCLIP 19.26 22.64 16.50 33.02 29.12 40.50 36.32 47.08 40.86 48.10 63.28 37.28 27.88 50.56 43.64 37.07
Ours 27.68 29.50 28.74 34.52 30.58 39.08 35.50 47.08 42.26 48.48 59.60 39.84 30.20 48.42 40.88 38.82
D Comparison to OT-VP
The very recent OT-VP [zhang2024ot] has presented a solution integrating optimal transport for test-time adaptation. Nevertheless, it presents several fundamental differences with our work. First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized. Thus, optimal transport is used for a different task. Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality. And third, the adaptation scenario strongly differs from our setting. In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome). While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 14. In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps. To expand the extent of our empirical validation, we compare our approach to OT-VP [zhang2024ot] whose results are reported in Table 14. To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
Table 14: Performance comparison to [zhang2024ot] across PACS and OfficeHome datasets. Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
Method PACS OfficeHome Avg.
First Setting : Single Source